OpenStudy (anonymous):

If h(x) = −2x − 10, find h(−4). −2 −18 −3 −16

OpenStudy (mayankdevnani):

h(-4)=-2(-4)-10=8-10=-2

OpenStudy (mayankdevnani):

ok @kristopher101495

OpenStudy (mayankdevnani):

option a

OpenStudy (anonymous):

ok thnx

OpenStudy (mayankdevnani):

welcome